# Systematic Constraint Formulation and Collision Free Trajectory Planning Using Space-Time Graphs of Convex Sets

### v1

@@ Abstract
In this paper, we create optimal, collision-free, time-dependent trajectories through cluttered dynamic environments. The many spatial and temporal constraints make finding an initial guess for a numerical solver difficult. Graphs of Convex Sets (GCS) and the recently developed Space-Time Graphs of Convex Sets formulation (ST-GCS) enable us to generate optimal minimum distance collision-free trajectories without providing an initial guess to the solver. We also explore the derivation of general GCS-compatible constraints and document an intuitive strategy for adapting general constraints to the framework. We show that ST-GCS produces results that are equivalent to the standard GCS formulation when the environment is static. We then show ST-GCS operating in dynamic environments to find minimum distance collision-free trajectories.
I Introduction
Figure 1: A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$ (a). Safe regions of space and time are partitioned into a Graph of Convex Sets (GCS), and then a collision-free trajectory is generated through the graph (b).
@@ I Introduction
Continuous optimization methods, by contrast, can directly enforce physical constraints on trajectories. These methods model obstacles using analytic functions such as signed distance fields [12], control barrier functions [13], or polygonal representations [14], and they optimize a cost function subject to these constraints. However, a key limitation of continuous optimization is its reliance on a good initial guess. In complex or cluttered environments, poor initializations can lead to suboptimal or infeasible solutions. To mitigate this, sampling-based methods are often used to generate feasible initial trajectories [15, 16]. Still, these approaches are especially sensitive to moving obstacles, which introduce additional temporal nonlinearities and local minima [17, 18].
@@
A limitation of this approach is that the current GCS framework requires that only one convex set in the graph may intersect the final trajectory state. As a result, Tang et al. conservatively select the final set, potentially preventing the framework from finding true minimum time trajectories even if they exist. Additionally, their formulation is restricted to planning piecewise-constant velocity trajectories and is not applied to higher-order splines. Critical constraints that ensure causality within the space–time configuration space are present, but their importance is not discussed or elaborated.
@@
This paper makes two main contributions. First, we provide a systematic methodology, accompanied by step-by-step examples, for rewriting common constraints in the form required by the GCS framework. Second, building on the ST-GCS framework, we investigate the generation of minimum-distance, collision-free trajectories in dynamic 2D environments. By assuming known obstacle positions and constant obstacle velocities, our approach produces higher-order optimal spline trajectories without relying on nonlinear solvers, initial temporal guesses, or incurring unnecessary computational complexity.
The remainder of this paper is structured as follows. Section II provides an overview of the underlying mathematics for graphs of convex sets and Bézier Splines. Section III details our methodology and presents examples of formulating GCS-compatible constraints. Section IV presents the results of optimizing trajectories with GCS in dynamic environments. Finally, Section V concludes the paper and discusses directions for future work.
II Background
The GCS optimization framework provides strong guarantees for the end user, and it is rooted in advanced concepts from convex analysis and optimization theory. In this section, we will highlight the key concepts for understanding GCS at a high level. For a more comprehensive understanding of GCS, we refer readers to [19, 20, 21, 22, 23, 24, 25, 26, 28].
II-A Convex Tools
@@
Figure 2: A nonconvex set $\mathcal{S}$ shown in (a) and its associated convex hull $\text{conv}(\mathcal{S})$ shown in (b).
@@
Figure 4: A set $\mathcal{S}\in\mathbb{R}^{2}$ (a) and the closure of the homogenization of $\mathcal{S}$ (b). The $\text{clh}(\mathcal{S})$ forms a cone in $\mathbb{R}^{3}$.
II-B Graph Theory
@@
II-C Graphs of Convex Sets (GCS) Optimization
@@
$\min_{x}\quad$ $\sum_{v\in\mathcal{V}}f_{v}(\vec{x}_{v})+\sum_{e\in\mathcal{E}}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (6a)
@@
$\min_{x,y}\quad$ $\sum_{v\in\mathcal{V}}y_{v}f_{v}(\vec{x}_{v})$
$+$ $\sum_{e\in\mathcal{E}}y_{e}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (7a)
@@
$\sum y_{(a,v)}+\delta_{s}$
$=$ $\sum y_{(v,b)}+\delta_{f}=y_{v}\quad\forall v\in\mathcal{V}$ (8e)
$\vec{z}_{v}\in y_{v}\mathcal{X}_{v}\quad\forall v\in\mathcal{V}$ (8f)
$\vec{z}_{e,a}\in y_{e}\mathcal{X}_{e}\quad\forall e\in\mathcal{E}$ (8g)
$\vec{z}_{e,b}\in y_{e}\mathcal{X}_{e}\quad\forall e\in\mathcal{E}.$ (8h)
@@
II-D Generating The Graph Of Convex Sets
@@ II-E Bézier Splines
Figure 7: Bézier curves associated with sets in GCS (a) and the solution to the GCS problem that connects Bézier curves to form a spline from start to end (b). Unused curves are removed.
III Methods
@@ III-A GCS Constraints
Our method for writing GCS-compatible edge constraints is summarized in Algorithm 1. The first step is to write the constraint as if it were active for a single edge in the solution subgraph. Then, both sides of the equation are multiplied by the decision variable that represents whether that edge is active. Next, the constraints of all edges flowing out of a single vertex are then summed together and simplified, resulting in a convex combination of potential constraints. Then, the bilinear products are replaced with the bilinear variable $z$. Finally, the constraint is repeated for each vertex in the graph, excluding the sink vertex.
@@
1:Write constraint $C$ for a single edge, in terms of the vertex control points $x_{v}$ and the connecting control points $\vec{x}_{b}$.
2:Multiply both sides of the constraint by the edge decision variable $y_{e}$.
3:Sum the constraints for all edges connected to the vertex $v$.
4:Simplify the expression using the relationship in Equation (8e) assuming that $\delta_{f}=0$, $y_{v}=\sum y_{(v,b)}$.
5:Substitute bilinear products with the variables $\vec{z}_{v}=y_{v}\vec{x}_{v}$ and $\vec{z}_{e,v}=y_{e}\vec{x}_{v}$.
6:Repeat the constraint for each vertex $v\in\mathcal{V},v\neq\text{sink}$.
@@ III-C Continuity Constraints
$\vec{x}_{v,n}=\vec{x}_{b,0}.$ (13)
Here, $\vec{x}_{v,n}$ denotes the last control point of the current curve in set $v$, and $\vec{x}_{b,0}$ denotes the first control point of the connecting set $b$.
We then multiply both sides of the equation by the edge decision variable connecting set $v$ with set $b$, $y_{(v,b)}$, making the constraint fully active when the edge is included in the solution and inactive otherwise:
$y_{(v,b)}\vec{x}_{v,n}=y_{(v,b)}\vec{x}_{b,0}.$ (14)
Next, we sum these constraints over all edges connected to vertex $v$:
$\sum_{b}y_{(v,b)}\vec{x}_{v,n}=\sum_{b}y_{(v,b)}\vec{x}_{b,0}.$ (15)
Assuming the flow conservation constraint from Equation 8e holds, the left-hand side simplifies to:
$y_{v}\vec{x}_{v,n}=\sum_{b}y_{(v,b)}\vec{x}_{b,0}.$ (16)
Figure 8: A vertex $v$ with adjacent vertices $b_{i}$. The continuity constraint ensures that the final Bézier control point falls in the convex hull of the connecting edges’ first Bézier control point (a), and then connects two edges as $y_{v,b_{i}}\rightarrow 1$ (b).
Referring back to Equations (3) and (4), this constraint ensures that the last control point of a set lies within the convex hull of the first control points of the connected sets, weighted by their corresponding decision variables $y_{(v,b)}$. When applied with the constraint requiring all control points to remain within their respective convex sets, it ensures that the control points meet at the boundary between sets when $y=1$, as that is the only place where both constraints can be met simultaneously. This enforces continuity between curve segments when edges are active and enables smooth transitions in the convex relaxation. If all edges and vertices are inactive, the constraint has no effect. Because this formulation relates a vertex to its outgoing edges, it is applied to every vertex except the sink vertex, which is implicitly captured through the other edge constraints in the graph. Alternatively, the constraint could be formulated using ingoing edges, in which case the source vertex would be excluded instead. Figure 8 illustrates this constraint.
We then substitute the bilinear variables, $\vec{z}_{v,n}=y_{v}\vec{x}_{v,n}$ and $\vec{z}_{(v,b),0}=y_{(v,b)}\vec{x}_{b,0}$, to rewrite the constraint as:
$\vec{z}_{v,n}$ $=\sum_{i=0}^{n}\vec{z}_{(v,b),0}.$ (17)
III-D Differentiability Constraints
@@
For the path to be differentiable where the two curves connect, the vector between the last two control points of one curve must equal the vector between the first two control points of the connecting curve. For the edge $(v,b)$, this constraint can be written as:
$\vec{x}_{v,n}-\vec{x}_{v,n-1}=\vec{x}_{b,1}-\vec{x}_{b,0}.$ (18)
As before, we multiply both sides of the equation by the edge decision variable $y_{(v,b)}$, making the constraint active when the edge is in the solution. The equation can be written as:
$y_{(v,b)}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (19)
We then sum these constraints over all edges connected to vertex $v$:
$\sum_{b}y_{(v,b)}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=\sum_{b}y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (20)
Assuming the flow conservation constraint from Equation 8e holds, the left-hand side simplifies to:
$y_{v}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=\sum_{b}y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (21)
We then substitute the bilinear variables, $\vec{z}_{v,0}=y_{v}\vec{x}_{v,n}$, $\vec{z}_{v,n-1}=y_{v}\vec{x}_{v,n-1}$, $\vec{z}_{(v,b),1}=y_{(v,b)}\vec{x}_{b,1}$, and $\vec{z}_{(v,b),0}=y_{(v,b)}\vec{x}_{b,0}$, to rewrite the constraint as:
$\vec{z}_{v,n}-\vec{z}_{v,n-1}$ $=\sum_{b}(\vec{z}_{(v,b),1}-\vec{z}_{(v,b),0}).$ (22)
However, this formulation is inefficient because it requires defining two $z$ variables for every directed edge, resulting in $|\mathcal{E}|\times 2$ extra variables to enforce this constraint, where $|\mathcal{E}|$ represents the cardinality of the edges of the graph. To reduce the number of variables, we introduce a final substitution variable, $\vec{z}_{(v,b_{i}),\text{diff}}=\vec{z}_{(v,b_{i}),1}-\vec{z}_{(v,b_{i})}$. This allows us to express the differentiability constraint as:
$(\vec{z}_{v,n}-\vec{z}_{v,n-1})=\sum_{i=0}^{n}\vec{z}_{(v,b_{i}),\text{diff}}.$ (23)
This formulation only requires $|\mathcal{E}|$ additional variables to enforce differentiability, which accelerates the optimization process.
III-E Space-Time Configuration Constraints
We now introduce the space-time configuration constraints, which are essential for generating realistic trajectories in the space-time configuration space. In the space-time configuration space, it is possible to construct splines that are non-causal and result in infinite velocities. To prevent this, we impose a constraint that will enforce causality and limit the maximum velocity on the spline. This constraint will be written in a way that is compatible with Equation 8f

### v2

@@ Abstract
In this paper, we create optimal, collision-free, time-dependent trajectories through cluttered dynamic environments. The many spatial and temporal constraints make finding an initial guess for a numerical solver difficult. Graphs of Convex Sets (GCS) and the recently developed Space-Time Graphs of Convex Sets (ST-GCS) enable us to generate minimum distance collision-free trajectories without providing an initial guess to the solver. We also explore the derivation of general GCS-compatible constraints and document an intuitive strategy for adapting general constraints to the framework. We show that ST-GCS produces equivalent trajectories to the standard GCS formulation when the environment is static, as well as globally optimal trajectories in cluttered dynamic environments.
keywords:
[…]
{highlights}
Repeatable strategy for writing and adapting edge constraints for the GCS framework
First fully worked examples of deriving vertex and edge constraints for the GCS framework, something that is nonintuitive and not shown in the literature
Exploration of space-time graphs of convex sets (ST-GCS) and how space-time specific constraints allow the GCS framework to produce globally optimal minimum-distance trajectories in dynamic environments
1 Introduction
Figure 1: (a) A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$. (b) Safe regions of space and time are partitioned into a Graph of Convex Sets (GCS), and then a collision-free trajectory is generated through the graph.
@@ 1 Introduction
Continuous optimization methods, by contrast, can directly enforce physical constraints on trajectories. These methods model obstacles using analytic functions such as signed distance fields [12], control barrier functions [13, 14], or polygonal representations [15], and they optimize a cost function subject to these constraints. However, a key limitation of continuous optimization is its reliance on a good initial guess. In complex or cluttered environments, poor initializations can lead to suboptimal or infeasible solutions. To mitigate this, sampling-based methods are often used to generate feasible initial trajectories [16, 17, 18, 19]. Still, these approaches are especially sensitive to moving obstacles, which introduce additional temporal nonlinearities and local minima [20, 21].
@@
A limitation is that Tang et al. place conservative constraints on the final timing of the trajectory to ensure obstacle avoidance, potentially preventing the framework from finding true minimum time trajectories even if they exist. Additionally, their formulation is restricted to planning piecewise-constant velocity trajectories and is not applied to higher-order splines. Critical constraints that ensure causality within the space–time configuration space are present, but their importance is not discussed or elaborated.
@@
This paper makes two main contributions. First, we provide a systematic methodology, accompanied by step-by-step examples, for rewriting common constraints in the form required by the GCS framework. Second, building on the ST-GCS framework, we investigate the generation of minimum-distance, collision-free trajectories in dynamic 2D environments. By assuming known obstacle positions and constant obstacle velocities, our approach produces time-dependent, higher-order optimal spline trajectories without relying on nonlinear solvers, initial temporal guesses, or incurring unnecessary computational complexity.
The remainder of this paper is structured as follows. Section 2 provides an overview of the underlying mathematics for graphs of convex sets and Bézier Splines. Section 3 details our methodology and presents examples of formulating GCS-compatible constraints. Section 4 presents the results of optimizing trajectories with GCS in dynamic environments. Finally, Section 5 concludes the paper and discusses directions for future work.
2 Background
The GCS optimization framework offers robust guarantees for the end user and is grounded in concepts from convex analysis and optimization theory. In this section, we will highlight the key concepts for understanding GCS at a high level. For a more comprehensive understanding of GCS, we refer readers to [22, 23, 24, 25, 26, 27, 28, 29, 31].
2.1 Convex Tools
@@
Figure 2: (a) A nonconvex set $\mathcal{S}$. (b) The associated convex hull, $\text{conv}(\mathcal{S})$.
@@
Figure 4: (a) A set $\mathcal{S}\in\mathbb{R}^{2}$. (b) the closure of the homogenization of $\mathcal{S}$. The $\text{clh}(\mathcal{S})$ forms a cone in $\mathbb{R}^{3}$.
2.2 Graph Theory
@@
2.3 Graphs of Convex Sets (GCS) Optimization
@@
$\min_{x}\qquad$ $\sum_{v\in\mathcal{V}}f_{v}(\vec{x}_{v})+\sum_{e\in\mathcal{E}}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (6a)
@@
$\min_{x,y}\qquad$ $\sum_{v\in\mathcal{V}}y_{v}f_{v}(\vec{x}_{v})+\sum_{e\in\mathcal{E}}y_{e}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (7a)
@@
$\sum y_{(a,v)}+\delta_{s}=\sum y_{(v,b)}+\delta_{f}=y_{v}\ \ \forall v\in\mathcal{V}$ (8e)
$\vec{z}_{v}\in y_{v}\mathcal{X}_{v}\qquad\forall v\in\mathcal{V}$ (8f)
$\vec{z}_{e,a}\in y_{e}\mathcal{X}_{e}\qquad\forall e\in\mathcal{E}$ (8g)
$\vec{z}_{e,b}\in y_{e}\mathcal{X}_{e}\qquad\forall e\in\mathcal{E}.$ (8h)
@@
2.4 Generating The Graph Of Convex Sets
@@ 2.5 Bézier Splines
Figure 7: (a) Bézier curves associated with sets in GCS and (b) the solution to the GCS problem that connects Bézier curves to form a spline from start to end. Unused curves are removed.
3 Methods
@@ 3.1 GCS Constraints
Our method for writing GCS-compatible edge constraints is summarized in Algorithm 1. The first step is to write the constraint as if it were active for a single edge in the solution subgraph. Then, both sides of the equation are multiplied by the associated edge variable. Next, the constraint is summed with respect to ingoing and outgoing edges and simplified using edge flow conservation. Then, the bilinear products for control points are replaced with the bilinear variable $z$. Slack variables are introduced to represent the relationship across the edge and constrained to turn off when the edge turns off. Finally, the constraint is repeated for each vertex in the graph.
@@
1:Write the constraint for a single edge, in terms of the control points within set $a$, $\vec{x}_{a}$, and the connecting control points within set $b$, $\vec{x}_{b}$.
2:Multiply both sides of the constraint by the edge decision variable $y_{(a,b)}$.
3:Sum the constraints for all outgoing edges connected to the vertex $a$ and separately sum the constraints for all ingoing edges connected to the vertex $b$.
4:Simplify the expression using the relationship in Equation (8e). Assume that $\delta_{f}=\delta_{s}=0$.
5:Substitute control point bilinear products, $y_{v}x_{v,i}$ with $z_{v,i}$ variables. Then introduce a substitution variable that represents the constraint relationship in step 2.
6:Simplify as needed.
7:Constrain the substitution variable as appropriate, ensuring that it turns off with respect to the edge decision variable $y_{(a,b)}$.
8:Repeat the constraint for each vertex $v\in\mathcal{V}$.
@@ 3.3 Continuity Constraints
$\vec{x}_{a,n}=\vec{x}_{b,0}.$ (13)
Here, $\vec{x}_{a,n}$ denotes the last control point of the curve in set $a$, and $\vec{x}_{b,0}$ denotes the first control point of a connecting set $b$.
We then multiply both sides of the equation by the edge decision variable connecting set $a$ with set $b$, $y_{(a,b)}$, making the constraint fully active when the edge is included in the solution and inactive otherwise:
$y_{(a,b)}\vec{x}_{a,n}=y_{(a,b)}\vec{x}_{b,0}.$ (14)
Next, we sum these constraints over all outgoing edges connected to vertex $a$ and all ingoing edges connected to vertex $b$:
$\sum_{b}y_{(a,b)}\vec{x}_{a,n}=\sum_{b}y_{(a,b)}\vec{x}_{b,0}$ (15)
$\sum_{a}y_{(a,b)}\vec{x}_{a,n}=\sum_{a}y_{(a,b)}\vec{x}_{b,0}.$ (16)
We write the constraints in both directions to force the constraint slack variables to be applied properly on both sides of the edge. With only one side being accounted for, the constraint will only apply to the initial or final control point within a set.
Assuming the flow conservation constraint from Equation 8e holds so that
$y_{a}=\sum_{b}y_{(a,b)}$ (17)
$y_{b}=\sum_{a}y_{(a,b)},$ (18)
the control points can be removed from the sums and Equation (15) and Equation (16) simplify to
$y_{a}\vec{x}_{a,n}=\sum_{b}y_{(a,b)}\vec{x}_{b,0}$ (19)
$y_{b}\vec{x}_{b,0}=\sum_{a}y_{(a,b)}\vec{x}_{a,n}.$ (20)
Figure 8: (a) A vertex $a$ with adjacent vertices $b_{i}$. The continuity constraint ensures that the final Bézier control point falls in the convex hull of the connecting edges’ first Bézier control point. (b) The two edges connect as $y_{(a,b_{2})}\rightarrow 1$.
Referring back to Equations (3) and (4), this constraint ensures that the last control point of a set lies within the convex hull of the first control points of the outgoing sets, and the first control point of a set lies within the convex hull of the last control points of the ingoing sets, weighted by their corresponding decision variables $y_{(a,b)}$. This enforces continuity between curve segments when edges are active, enabling smooth transitions in the convex relaxation. Figure 8 illustrates Equation (19).
We then substitute the bilinear variables, $\vec{z}_{a,n}=y_{a}\vec{x}_{a,n}$, $\vec{z}_{b,0}=y_{b}\vec{x}_{b,0}$, to put the constraint in terms of the variables which represent the Bèzier control points. From Equation (14) we substitute $\vec{p}_{(a,b)}=y_{(a,b)}\vec{x}_{a,n}=y_{(a,b)}\vec{x}_{b,0}$ to represent the point where both spline segments meet, and rewrite the constraint as:
$\vec{z}_{a,n}$ $=\sum_{b}\vec{p}_{(a,b)},$ (21)
$\vec{z}_{b,0}$ $=\sum_{a}\vec{p}_{(a,b)}.$ (22)
Finally, we constrain the point where the splines meet to lie on the intersection between the two convex sets, coupled by the edge decision variable $y_{(a,b)}$:
$\begin{bmatrix}A_{a}\\ A_{b}\end{bmatrix}\vec{p}_{(a,b)}\;\leq\;\begin{bmatrix}\vec{d}_{a}\\ \vec{d}_{b}\end{bmatrix}y_{(a,b)}.$ (23)
3.4 Differentiability Constraints
@@
For the path to be differentiable where the two curves connect, the vector between the last two control points of one curve must equal the vector between the first two control points of the connecting curve. For the edge $(a,b)$, this constraint can be written as:
$\vec{x}_{a,n}-\vec{x}_{a,n-1}=\vec{x}_{b,1}-\vec{x}_{b,0}.$ (24)
As before, we multiply both sides of the equation by the edge decision variable $y_{(a,b)}$, making the constraint active when the edge is in the solution. The equation can be written as:
$y_{(a,b)}(\vec{x}_{a,n}-\vec{x}_{a,n-1})=y_{(a,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (25)
Next, we sum these constraints over all outgoing edges connected to vertex $a$ and all ingoing edges connected to vertex $b$:
$\sum_{b}y_{(a,b)}(\vec{x}_{a,n}-\vec{x}_{a,n-1})=\sum_{b}y_{(a,b)}(\vec{x}_{b,1}-\vec{x}_{b,0})$ (26)
$\sum_{a}y_{(a,b)}(\vec{x}_{a,n}-\vec{x}_{a,n-1})=\sum_{a}y_{(a,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (27)
Assuming the flow conservation constraint from Equation 8e holds, the equations can be simplified and rearranged to
$y_{a}(\vec{x}_{a,n}-\vec{x}_{a,n-1})$ $=\sum_{b}y_{(a,b)}(\vec{x}_{b,1}-\vec{x}_{b,0})$ (28)
$y_{b}(\vec{x}_{b,1}-\vec{x}_{b,0})$ $=\sum_{a}y_{(a,b)}(\vec{x}_{a,n}-\vec{x}_{a,n-1}).$ (29)
Then, we distribute the $y$ variables through the left-hand side and substitute the bilinear variables, $\vec{z}_{a,n}=y_{a}\vec{x}_{a,n}$, $\vec{z}_{a,n-1}=y_{a}\vec{x}_{a,n-1}$, $\vec{z}_{b,1}=y_{b}\vec{x}_{b,1}$, $\vec{z}_{b,0}=y_{b}\vec{x}_{b,0}$. After this, we use Equation (25) and substitute $\vec{q}_{(a,b)}=y_{(a,b)}(\vec{x}_{a,n}-\vec{x}_{a,n-1})=y_{(a,b)}(\vec{x}_{b,1}-\vec{x}_{b,0})$ to represent the difference between control points. The constraint can be written as:
$\vec{z}_{a,n-1}$ $=\vec{z}_{a,n}-\sum_{b}\vec{q}_{(a,b)}$ (30)
$\vec{z}_{b,1}$ $=\vec{z}_{b,0}+\sum_{a}\vec{q}_{(a,b)}.$ (31)
Finally, we constrain the 2-norm of $\vec{q}_{(a,b)}$ by some finite constant value $\gamma$, coupled with the edge decision variable,
$\left\lVert\vec{q}_{(a,b)}\right\rVert_{2}\leq\gamma y_{(a,b)}.$ (32)
3.5 Space-Time Configuration Constraints
We now introduce the space-time configuration constraints, which are essential for generating realistic trajectories in the space-time configuration space. In the space-time configuration space, it is possible to construct splines that are non-causal and result in infinite velocities. This happens when the time component of the trajectory is not unique, meaning that for certain values of $t$ there are multiple corresponding locations on the trajectory. To prevent this, we impose a constraint that will enforce causality and limit the maximum velocity on the spline. This constraint will be written in a way that is compatible with Equation 8f